OpenStudy (anonymous):

A skier moves 100.0 m horizontally, and then travels another 35.0 m uphill at an angle of 35.0º above the horizontal. What is the skier's displacement from his starting point?

OpenStudy (anonymous):

Displacement is the minimum distance between two points.  Calculate AB

OpenStudy (anonymous):

You can use "the law of parallelograms". \[AB = \sqrt{AC ^{2}+CB ^{2} - 2 \times AC \times CB \cos <ACB }\] angle ACB is 180 - 35 = 145 deg